# ManiFoundation Model for General-Purpose Robotic Manipulation of Contact Synthesis with Arbitrary Objects and Robots

### v1

@@ IV Dataset
We create a large-scale comprehensive annotated dataset that includes articulated/rigid objects and deformable objects in 1D/2D/3D forms. See Appendix B-A, B-B, B-C, and our website for more details.
@@
To systematically evaluate the model performance, we design $2\times 3=6$ test sets. The “3” here stands for three types of objects (rigid, cloth, 3D deformable). The “2” here is for two test groups: Seen Object & Unseen Motion and Unseen Object & Unseen Motion. Specifically, the Seen Object & Unseen Motion test group is designed to assess whether our model can learn to manipulate an object to reach an unseen arbitrary goal after seeing the object in the training set. The Unseen Object & Unseen Motion test group is designed to evaluate our model’s generalization to novel objects. All 6 test sets contain $\sim$200 examples.
@@ V Experiments
  • Q3: How do physical properties affect the output of the network?
@@
  • Q5: Can manipulation region mask affect and constrain the network contact heatmaps and the optimization result?
  • Q6: How effective is our optimization algorithm compared to other approaches on force-closure grasp optimization?
  • Q7: Does combining network and optimization improve task-oriented grasp generation? (See Supplementary)
  • Q8: Can our model generalize to novel robot manipulators? (See Supplementary)
  • Q9: How does our proposed framework perform on real robot experiments with real point clouds? (See Supplementary)
@@
For deformable objects: We apply the predicted contacts on the objects in simulation. We use DiffCloth [26] for 2D deformable objects and SoftMAC [27] for 1D/3D deformable objects. We can calculate the object point cloud after applying contact forces in simulators. The success rate is defined based on the distance between the results from the object point cloud and the target point cloud. Detailed thresholds and parameters are put in the supplementary.
@@ V-A1 Overall Performance & Data Fusion
Figure 4: Visualizations of evaluation on deformable objects. The color of the point cloud represents the contact heatmap, with purple representing low values and yellow representing high values. Arrows are the force or motion direction. The representation of all visualizations of this figure is consistent with the latter figures.
To answer Q1 and Q2, we train our model described in Sec. III-A on the above 3K training set denoted as Ours(3K) and compare it with three approaches: Baseline, Ours(SD), and Ours. Baseline randomly selects the same number of contact points as Ours, applying forces or motions in random directions for deformable objects. For articulated/rigid object, Baseline sets the robot hand/fingers at the maximum open status, then moves it toward the object’s center, closing the fingers upon contact with the object’s surface. Baseline calculate the forces of each fingerstip using Eqn. 7. We chose such a baseline because we are the first to train a unified model for contact synthesis on objects of various types. We put other less related baselines in supplementary. Ours(SD) stands for training our model only Single Dataset ($\sim$1K rigid, $\sim$1K clothes, $\sim$1K 3D deformation respectively instead of together). Ours contains our network trained on a larger dataset of $\sim$200K training data.
To answer Q1, as shown in Tab. I, Ours has an average success rate of over 90, 85, 80% over rigid objects, 3D deformable objects, and 2D deformable objects. It indicates our pipeline can synthesize reasonable contacts for diverse objects with broad generalization. Meanwhile, the low success rate of Baseline and the visualization 4 illustrates the difficulty of each task.
To address Question 2, we conducted a comparison between Ours(3K) and Ours(SD). The Ours(3K) have comparable or slightly better performance. This reflects that our model is capable of effectively integrating various object types without any degradation in performance, compared to training exclusively on a single object type.
@@ V-A2 Physical Properties
Figure 5: Visualizations of how physical properties affect network outputs. Blue squares and cylinders are object meshes. Overlapping images with transparency represent the moved object. The orange arrow represents the motion of the object.
To further understand the impact of the physical properties (in Q3), we conducted two experiments depicted in Fig. 5, evaluating how variations in physical properties influence our model’s behavior. The experiments involved two shapes: a thin square and a long cylinder. The square, when associated with the physical properties of cloth or a rigid body, represents a deformable handkerchief or a rigid card, respectively. Similarly, the cylinder, with the properties of a 3D deformable object or a rigid body, symbolizes a rope or a stick. We use identical point clouds and motions to our model, altering only the physical properties, and observing the outcomes.
Fig. 5 demonstrates that to advance the thin square, our model opts to pull the forward corners, reflecting the handkerchief’s pliability, and pushes from the rear for the rigid card. Pushing the handkerchief from behind would cause significant deformation. For the long cylinder, our model drags both ends on the rope and pushes the middle point given stick’s rigidity. The above result illustrates that our model can adjust its outputs based on the input of physical properties, indicating our model learns to reason about the dynamics of the object.
@@ V-A3 Multimodal Syntheses
To answer Q4 - whether our model can generate diverse solutions given the same input, we choose four novel objects that are not in our training dataset. For each object, we choose a common desired object motion such as rotating the knob and tilting the book. Given the above input, we draw different $z\in\mathbb{R}^{64}$ samples from $\mathcal{N}(0,I)$ and visualize their outputs. As shown in Fig. 6, thanks to our CVAE module design, our method can generate diverse physically plausible contact outputs given the same input. For example, to rotate a knob, our model can provide solutions that include both holding the upper side and manipulating the lower side of the knob.
@@ V-A4 Manipulation Region Selection
To explore whether a manipulation region mask can effectively guide the network’s contact point predictions (Q5), we tested our model with three new everyday objects and tasks, not included in our training set. These objects have specific manipulation areas, like handles. We assessed the model’s performance with and without a manually selected manipulation region mask, and the results are displayed in Fig. 7. The mask leads the model to favor these regions. Without a mask, it may suggest contact points suitable for the intended motion but unsuitable for practical manipulation. For instance, lifting a cup from its bottom on a table is unfeasible. Applying a mask to the handle helps the model to propose viable contact points, preventing potential collisions. This approach’s effectiveness is depicted in the robot hand poses generated by our model, shown on the right side of Fig. 7. This method’s success is illustrated by our model’s generated hand poses, presented in Fig. 7. Although generating the ideal manipulation mask isn’t our focus, our model includes an API for users to define their manipulation regions.
@@ V-B Evaluating Robot Hand Pose Optimization
We evaluate the effectiveness of our proposed robotic hand pose refinement for task-oriented grasp and force-closure grasp. Task-oriented grasp requires the grasp pose to produce a specific target wrench. Note that we have derived a relationship between the target wrench and target motion for rigid objects in Sec. III-B3. For simplicity in this evaluation, we use the target wrench as the manipulation task description. Our evaluation encompasses both task-oriented grasp and force-closure grasp scenarios over 1000 objects in our dataset. For each object, 20 distinct grasps are generated. Together, 20k task-oriented grasp and 20k force-closure grasp. In task-oriented evaluations, the target wrench for each grasp is randomly sampled. For this experiment, a grasp is considered successful if it meets three criteria: the residual wrench is less than $10^{-6}$, the distance between the fingertip and the object is less than 5mm, and there are no intersections between fingers or between fingers and the object.
For a fair comparison, we provide every method with the same initialization. We retrieve a ground truth 6D palm pose that can exert the target wrench or in force-closure from our dataset. We put this 6D palm pose 30cm away from the object and add a random 60-degree rotation and 5 cm translation to the ground truth palm pose and set all other finger joints at rest joint values. Under this initialization, we run our robotic hand pose refinement and compare it with a direct grasp method. The direct grasp method directly moves the robot hand toward the object’s center, closing the fingers upon contact with the object’s surface and calculating the forces of each fingertip using Eqn. 7.
Furthermore, our analysis includes three types of robotic hands: Barrett [33], LeapHand [34], and ShadowHand [35]. The results are shown in Table. II. Our refinement module provides much higher success rates than the baseline for both scenarios and all considered hands.
@@ V-C Real World Experiments
In this section, we evaluate our system’s performance of manipulating various rigid, clothes, and 3D deformation objects in multiple real-world settings, as shown in Figure 1. We put detailed descriptions in the supplementary. Rope Rearrangement We design a rope rearrangement experiment for a Kinova MOVO robot arm to rearrange a random reset rope to letters “NUS”. We perform 10 times for each letter in “NUS”. Our model achieves 90% success rate, indicates that our model is accurate in generating grasp points for deformation objects like ropes. Breakfast Preparation We design a complex breakfast preparation experiment for a single Flexiv robot arm and a LeapHand as the manipulator. In this setting, the robot will open the fridge door, take out the milk box, place the milk box on the table, pick up a piece of bread, put the bread into the toaster, open the toaster, take out the cooked bread and place it to the plate. Among these objects, the fridge door is an articulated object, while the milk box is a normal rigid body object, and the bread is treated as a 3D deformation object. Cloth Folding We design a cloth folding experiment for a Kinova MOVO robot arm to fold a T-shirt. We extract the T-shirt point clouds using SAM [36]. Movo will then grasp the point with the highest heatmap value and move to the predicted target point.
@@ VI Conclusion and Limitations
We introduce a framework that develops a manipulation foundation model via contact synthesis, taking inputs like point clouds, physical properties, target motions, and region masks to output contact points and forces or motions. Our model, tested on a variety of objects including articulated, rigid, and deformable items, demonstrates average success rates of around 90% across simulations and real-world applications.
Our model can be further improved in the following aspects. First, our framework focuses on quasi-static tasks because we separate the whole trajectory into discrete motions at each time step. Future works can seek to extend our framework to high-dynamical tasks by considering multiple-step motions or the whole trajectory at each timestep. Secondly, our model focuses on single-hand finger-point contact. Future works can seek to extend our framework to multiple hand surface contact situations.

### v2

@@ IV Dataset
We create a large-scale comprehensive annotated dataset that includes articulated/rigid objects and deformable objects in 1D/2D/3D forms. Refer to our  website for more details.
@@
To systematically evaluate the performance, we design $2\times 3=6$ test sets. The “3” here stands for three types of objects (rigid, cloth, 3D deformable). The “2” here is for two test groups: Seen Object & Unseen Motion and Unseen Object & Unseen Motion. Specifically, the Seen Object & Unseen Motion test group is designed to assess whether our model can learn to manipulate an object to reach an unseen arbitrary goal after seeing the object in the training set. The Unseen Object & Unseen Motion test group is designed to evaluate our model’s generalization to novel objects. Each test set contains $\sim$200 examples.
@@ V Experiments
  • Q3: How do physical properties affect the model output?
@@
  • Q5: Can manipulation region mask constrain the network contact heatmaps and the optimization result?
  • Q6: How effective is our grasp optimization?
  • Q7: Does combining network and optimization improve task-oriented grasp generation? (See Supp.)
  • Q8: Can our model generalize to novel robot manipulators? (See Supp.)
  • Q9: How does our proposed framework perform on real robot experiments with real point clouds? (See Supp.)
@@
For deformable objects: We apply the predicted contacts on the objects in simulation. We use DiffCloth [24] for 2D deformable objects and SoftMAC [25] for 1D/3D deformable objects. We apply contact forces to the object and obtain the rollout point cloud in simulators. The success is determined by the distance between the rollout point cloud and the target point cloud. Detailed thresholds and parameters are in Supp.
@@ V-A1 Overall Performance & Data Fusion
Figure 4: Visualizations of evaluations on deformable objects are shown. The point cloud colors represent the contact heatmap, with purple indicating low values and yellow indicating high values. Arrows denote the force or motion direction. This representation is consistent across all figures.
To answer Q1 and Q2, we train our model described in Sec. III-A on the above 3K training set denoted as Ours(3K) and compare it with three approaches: Baseline, Ours(SD), and Ours. Baseline randomly selects the same number of contact points as Ours, applying forces or motions in random directions for deformable objects. For articulated/rigid object, Baseline sets the robot hand/fingers at the maximum open status, then moves it toward the object’s center, closing the fingers upon contact with the object’s surface. Baseline calculate the forces of each fingerstip using Eqn. 6. We chose such a baseline because we are the first to train a unified model for contact synthesis on objects of various types. Ours(SD) stands for training our model only on Single Dataset ($\sim$1K rigid, $\sim$1K clothes, $\sim$1K 3D deformation respectively instead of together). Ours is for our network trained on a $\sim$200K dataset training data.
To answer Q1, as shown in Tab. I, Ours has an average success rate of over 90, 85, 80% over rigid, 3D deformable, and 2D deformable objects. It indicates our pipeline can synthesize reasonable contacts for diverse objects with broad generalization. Meanwhile, the low success rate of Baseline and the visualization 4 illustrate the difficulty of each task.
To address Question 2, we conducted a comparison between Ours(3K) and Ours(SD). The Ours(3K) have comparable or slightly better performance. This reflects that our model is capable of effectively integrating various object types without performance loss, compared to training on a single object type.
@@ V-A2 Physical Properties
Figure 5: Visualizations of how physical properties affect network outputs. Blue squares and cylinders are objects. Overlapping images with transparency represent the moved object. The orange arrow represents the motion.
To understand the impact of physical properties (in Q3), we conducted two experiments shown in Fig. 5, evaluating how variations in physical properties influence our model’s behavior. We used two shapes: a thin square and a long cylinder. The square represented a handkerchief (deformable) or a card (rigid). The cylinder represented a rope (deformable) or a stick (rigid). We used identical point clouds and motions, altering only the physical properties to observe the outcomes. Fig. 5 shows that for the thin square, our model pulls the forward corners for the handkerchief, reflecting its pliability, and pushes from the rear for the rigid card. For the long cylinder, our model drags both ends of the rope and pushes the middle of the stick. These results demonstrate that our model adjusts its outputs based on the physical properties, indicating it learns to reason about object dynamics.
@@ V-A3 Multimodal Syntheses
To answer Q4—whether our model can generate diverse solutions from the same input—we tested four novel objects not in our training dataset. For each object, we selected a common motion, such as rotating a knob or tilting a book. Using these inputs, we sampled different $z\in\mathbb{R}^{64}$ values from $\mathcal{N}(0,I)$ and visualized the outputs. As shown in Fig. 6, our CVAE module allows our model to generate diverse, physically plausible contact outputs. For example, to rotate a knob, the model provides solutions like holding the upper or the lower side.
@@ V-A4 Manipulation Region Selection
To explore whether a manipulation region mask can guide the network’s contact point predictions (Q5), we tested our model with three new everyday objects and tasks not included in our training set. These objects have specific manipulation areas, like handles. We assessed the model’s performance with and without a manually selected manipulation region mask, and the results are shown in Fig. 7. The mask guides the model to favor these regions. Without a mask, it may suggest contact points suitable for the intended motion but impractical for manipulation. For instance, lifting a cup from its bottom on a table is unfeasible. Applying a mask to the handle helps the model propose viable contact points, preventing potential collisions. The effectiveness of this approach is illustrated by the robot hand poses generated by our model, shown on the right side of Fig. 7. Although generating the ideal manipulation mask isn’t our focus, our model includes an API for users to define their manipulation regions.
@@ V-B Evaluating Robot Hand Pose Optimization
We evaluate the effectiveness of our proposed robotic hand pose refinement for task-oriented and force-closure grasps. Task-oriented grasp requires the grasp pose to produce a specific target wrench. As derived in Sec. III-B3, we use the target wrench as the manipulation task description for rigid objects. Our evaluation includes both task-oriented and force-closure grasp scenarios over 1000 objects in our dataset, generating 20 distinct grasps per object, totaling 20k task-oriented and 20k force-closure grasps. In task-oriented evaluations, the target wrench for each grasp is randomly sampled. A grasp is considered successful if it meets three criteria: the residual wrench is less than $10^{-6}$, the distance between the fingertip and the object is less than 5mm, and there are no intersections between fingers or between fingers and the object.
For a fair comparison, we initialize each method identically. We retrieve a ground truth 6D palm pose from our dataset that can exert the target wrench or achieve force-closure. This 6D palm pose is placed 30cm away from the object, and then a random 60-degree rotation and 5cm translation are added, with all finger joints set at rest positions. Under this initialization, we run our robotic hand pose refinement and compare it with a direct grasp method. The direct grasp method moves the robot hand directly towards the object’s center, closes the fingers upon contact, and calculates the forces for each fingertip using Eqn. 6.
Our analysis includes three types of robotic hands: Barrett [31], LeapHand [32], and ShadowHand [33]. As shown in Table II, our refinement module achieves significantly higher success rates than the baseline for both scenarios and all tested hands.
@@ V-C Real World Experiments
In this section, we evaluate our system’s performance in manipulating various rigid objects, clothes, and 3D deformable objects in real-world settings, as shown in Fig. 1. Detailed descriptions are provided in the Supp. Rope Rearrangement We designed an experiment for a Kinova MOVO robot arm to rearrange a randomly reset rope into the letters ”NUS.” We performed the task 10 times for each letter, achieving a 90% success rate, indicating our model’s accuracy in generating grasp points for deformable objects like ropes. Breakfast Preparation We designed a complex breakfast preparation experiment using a single Flexiv robot arm with a LeapHand manipulator. The tasks included opening the fridge door (articulated object), taking out the milk box (rigid object), placing the milk box on the table, picking up a piece of bread, putting it into the toaster, opening the toaster, and placing the cooked bread on a plate. The bread is treated as a 3D deformable object. Cloth Folding We designed a cloth folding experiment for a Kinova MOVO robot arm to fold a T-shirt. The robot grasps the point with the highest heatmap value and moves to the predicted target point.
@@ VI Conclusion and Limitations
We introduce a framework for developing a manipulation foundation model for contact synthesis. This model takes inputs like point clouds, physical properties, target motions, and region masks to output contact points and forces or motions. Tested on various objects, including articulated, rigid, and deformable items, our model achieves average success rates of around 90% in both simulations and real-world settings.
Our model can be improved in the following aspects. First, our framework focuses on quasi-static tasks by separating the trajectory into discrete motions at each time step. Future work could extend the framework to high-dynamic tasks by considering multi-step motions or the entire trajectory at each timestep. Second, our model focuses on single-hand finger-point contact. Future work could extend the framework to multi-hand surface contact situations. Thirdly, our current model has about 5 million parameters. In the future, we plan to scale up by increasing the amount of data and the number of network parameters.